Dld

A Dynamic Link/Unlink EditorCopyright \begin{matrix}c \\ ◯ \end{matrix} 1990 W. Wilson Ho.by W. Wilson Ho

1. What is dld?

Dld is a library package of C functions that performs dynamic link editing. Programs that use dld can add compiled object code to or remove such code from a process anytime during its execution. Loading modules, searching libraries, resolving external references, and allocating storage for global and static data structures are all performed at run time.

Dld is now available for VAX, Sun 3, and SPARCstation machines.

This text describes how the dld functions can be called and some technical details that should be aware of. For the internals of dld and sample applications, please refer to the technical report An Approach to Genuine Dynamic Linking, CSE-90-25, Division of Computer Science, University of California at Davis, CA 95616 (email: how@ivy.ucdavis.edu).

1.1 Initializing Dld

To use any of the dld functions, you must include the header file dld.h for declaration of the functions and definition of the error code constants.

The function dld_init must be called before any other dld functions. It has the following syntax:

     int dld_init (char *path)

where path is a string containing the path name of the executable file of the executing process.

This function initializes internal data structures of dld and loads into memory symbol definitions of the executing process. By doing so, other dynamically loaded functions can reference symbols already defined or share functions already exist in the executing process.

dld_init returns 0 when successful; otherwise, it returns an error code that is non-zero (see <undefined> [Definition of Error Codes], page <undefined>).

1.1.1 Locating the Executable File

The path name of the executing process as required by dld_init might not be easily obtained all the time. Not all systems passes the entire path name of the executable file as the first argument (argv[0]) to main. In order to obtain the full path of the executable file, the dld_find_executable function can be used. This function has the following syntax:

     char *dld_find_executable (char *command)

dld_find_executable returns the absolute path name of the file that would be executed if command were given as a command. It looks up the environment variable PATH, searches in each of the directory listed for command, and returns the absolute path name for the first occurrence. Thus, it is advisable to invoke dld_init as:

     main (int argc, char **argv)
     {

…

         if (dld_init (dld_find_executable (argv[0]))) {

…

         }

…

     }

Note: If the current process is executed using the execve call without passing the correct path name as argument 0, dld_find_executable (argv[0]) will also fail to locate the executable file.

dld_find_executable returns zero if command is not found in any of the directories listed in

PATH.

1.2 Dynamically Linking in New Modules

The function dld_link dynamically links in the named relocatable object or library file into memory. It has the following syntax:

     int dld_link (char *filename)

where filename is the path name of the file to be linked. Specifically, if the named file is a relocatable object file, it is completely loaded into memory. If it is a library file, only those modules defining an unresolved external reference are loaded. Since a module in the library may itself reference other routines in the library, loading it may generate more unresolved external references. Therefore, a library file is searched repeatedly until a scan through all library members is made without having to load any new modules.

Storage for the text and data of the dynamically linked modules is allocated using malloc. In other words, they are kept in the heap of the executing process.

After all modules are loaded, dld_link resolves as many external references as possible. Note that some symbols might still be undefined at this stage, because the modules defining them have not yet been loaded.

If the specified module is linked successfully, dld_link returns 0; otherwise, it returns a non-zero error code (see <undefined> [Definition of Error Codes], page <undefined>).

1.3 Unlinking a Module

The major difference between dld and other dynamic linker is that dld allows object modules to be removed from the process anytime during execution. Unlinking a module is simply the reverse of the link operation (see <undefined> [Important Points in Using Unlink], page <undefined>). The specified module is removed and the memory allocated to it is reclaimed. Additionally, resolution of external references must be undone.

There are two unlink functions:

     int dld_unlink_by_file (char *path, int hard)
     int dld_unlink_by_symbol (char *id, int hard)

The two unlink functions are basically the same except that dld_unlink_by_file takes as argument the path name (path) of a file corresponding to a module previously linked in by dld_link, but dld_unlink_by_symbol unlinks the module that defines the specified symbol (id).

Both functions take a second argument hard. When hard is non-zero (hard unlink), the specified module is removed from memory unconditionally. On the other hand, if hard is zero (soft unlink), this module is removed from memory only if it is not referenced by any other modules. Furthermore, if unlinking a module results in leaving some other modules being unreferenced, these unreferenced modules are also removed.

Hard unlink is usually used when you want to explicitly remove a module and probably replace it by a different module with the same name. For example, you may want to replace the system’s printf by your own version. When you link in your version of printf, dld will automatically redirect all references to printf to the new version.

Soft unlink should be used when you are not sure if the specified module is still needed. If you just want to clean up unnecessary functions, it is always safe to use soft unlink.

Both unlink functions returns 0 if the specified object file or symbol is previously loaded. Otherwise, they return a non-zero error code (see <undefined> [Definition of Error Codes], page <undefined>).

1.3.1 Important Points in Using Unlink

When a module is being unlinked, dld tries to clean up as much as it can to restore the executing process to a state as if this module has never been linked. This clean up includes removing and reclaiming the memory for storing the text and data segment of the module, and un-defining any global symbols defined by this module.

However, side effects—such as modification of global variables, input/output operations, and allocations of new memory blocks—caused by the execution of any function in this module are not reversed. Thus, it is the responsibility of the programmer to explicitly carry out all necessary clean up operations before unlinking a module.

1.4 Invoking Dynamically Linked Functions

Dynamically linked functions may still be invoked from modules (e.g., main) that do not contain references to such functions. The function dld_get_func returns the entry point of the named function. This returned value can later be used as a pointer to the function. Similarly, the address of a global variable can be obtained by the function dld_get_symbol:

     unsigned long dld_get_symbol (char *id)
     unsigned long dld_get_func (char *func)

A typical use of dld_get_func would be:

     {
         void (*func) ();
         int error_code;

…

         /* First, link in the object file "my_object_file.o".
            Proceed only if the link operation is successful, i.e. it returns 0.
            "my_new_func" is a function defined in "my_object_file.o".
            Set func to point at the entry point of this function and then
            Invoke it indirectly through func. */
         if ((error_code = dld_link ("my_object_file.o")) == 0) {
             if ((func = (void (*) ()) get_func ("my_new_func")) != 0)
                 (*func) ();

…

         } else {

…

         }
     }

Both dld_get_func and dld_get_symbol return zero if the named function or symbol cannot be found.

1.5 Determining If a Function is Executable

Since dld allows modules to be added to or removed from an executing process dynamically, some global symbols may not be defined. As a result, an invocation of a function might reference an undefined symbol. We say that a function is executable if and only if all its external references have been fully resolved and all functions that it might call are executable.

The predicate function dld_function_executable_p helps solve this problem by tracing the cross references between modules and returns non-zero only if the named function is executable. It has the following syntax:

     int dld_function_executable_p (char *func)

Note that the implementation of dld_function_executable_p is not complete according to the (recursive) definition of executability. External references through pointers are not traced. That is, dld_function_executable_p will still return non-zero if the named function uses a pointer to indirectly call another function which has already been unlinked. Furthermore, if one external reference of a object module is unresolved, all functions defined in this module are considered unexecutable. Therefore, dld_function_executable_p is usually too conservative.

However, it is advisable to use dld_function_executable_p to check if a function is executable before its invocation. In such a dynamic environment where object modules are being added and removed, a function that is executable at one point in time might not be executable at another. Under most circumstances, dld_function_exectable_p is accurate. Also, the implementation of this function has been optimized and it is relatively cheap to use.

1.6 Definition of Error Codes

The dld functions return a non-zero error code when they fail. The definitions of these error codes are:

DLD_ENOFILE cannot open file. DLD_EBADMAGIC bad magic number. DLD_EBADHEADER failure reading header. DLD_ENOTEXT premature eof in text section. DLD_ENOSYMBOLS premature eof in symbols. DLD_ENOSTRINGS bad string table. DLD_ENOTXTRELOC premature eof in text relocation. DLD_ENODATA premature eof in data section. DLD_ENODATRELOC premature eof in data relocation. DLD_EMULTDEFS multiple definitions of symbol. DLD_EBADLIBRARY malformed library archive. DLD_EBADCOMMON common block not supported. DLD_EBADOBJECT malformed input file (not object file or archive). DLD_EBADRELOC bad relocation info. DLD_ENOMEMORY virtual memory exhausted. DLD_EUNDEFSYM undefined symbol.